Find derivatives of the following functions:

• y=(4x+3)/ln[2x2 +3x−5] • y=3(e2x +e−2x)(x2 −4)


1
Expert's answer
2022-01-17T09:54:09-0500

"y=\frac{4x+3}{\ln(2x^2+3x-5)},"

"y'=\frac{4\ln(2x^2+3x-5)-\frac{(4x+3)^2}{2x^2+3x-5}}{\ln^2(2x^2+3x-5)}."


"y=3(e^{2x}+e^{-2x})(x^2-4),"

"y'=6e^{-2x}(-x^2+e^{4x}(x^2+x-4)+x+4)."